OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form.

OpenStudy (anonymous):

Change into Slope-Intercept form. y = mx + b y - 2 = 2(x - 3) Distribute. y - 2 = 2x - 6 Add 2. y = 2x - 4 NOW change to standard form. Ax + By = C y = 2x - 4 Subtract 2x. -2x + y = -4 A can NOT be a negative. Multiply everything by -1. 2x - y = 4 <---Standard form

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome. (: